OpenStudy (anonymous):

determine whether the dot product is distributive with respect to the cross product.

OpenStudy (turingtest):

we want to know if\[v\cdot (u\times w)=(v\cdot u)\times (v\cdot w)\]\[???\]so write out the formula for each side and compare the results

OpenStudy (anonymous):

i think that's what it's asking and i did that and determined that it is not distributive

OpenStudy (turingtest):

correct

OpenStudy (anonymous):

but i honestly have no clue what i'm doing lol

OpenStudy (turingtest):

you may be able to get away with doing it with numbers, but they want you to set up random vectors and try to simplify I believe what part is troubling you?

OpenStudy (anonymous):

well i just wasn't sure if that's correct. i did use letters and i'm going to go with my answer. the next question is determine whether the cross product is associative. would i go u x (v x w) = (u x v) x w then determine if that does indeed equal each other or not

OpenStudy (turingtest):

exactly

OpenStudy (anonymous):

thanks !

OpenStudy (turingtest):

no prob, looks like you just needed confirmation :D